# Size $\neq$ Fairness: A Comprehensive Social Bias Audit of Contrastive Vision–Language Models

### v3

@@ Abstract
Size $\neq$ Fairness: A Comprehensive Social Bias Audit of Contrastive Vision–Language Models
@@ 1 Introduction
Existing audits seldom answer this question because they vary only one factor or inspect a single model family, leaving the effects of model capacity, data volume, and data composition entangled. In language-only settings, scaling can either help or hurt fairness depending on objective and metric (Liang et al., 2022), but systematic evidence for VLMs is still limited.
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani and de Melo, 2023), whereas DebiasCLIP prepends learned visual tokens to attenuate gender stereotypes (Berg and colleagues, 2022). Embedding-space analyses reveal that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2022). Prompt-level interventions, such as the Debiasing and VisDebiasing prefixes in ModSCAN, reduce but do not eradicate skew (Jiang et al., 2024). Beyond CLIP, BendVLM performs nonlinear, test-time debiasing of VLM embeddings without fine-tuning (Gerych et al., 2024), Association-Free Diffusion mitigates object-people stereotype transfer in text-to-image generation (Zhou et al., 2024), FairCoT leverages chain-of-thought reasoning in multimodal LLMs to iteratively refine prompts and improve fairness in text-to-image generation (Al Sahili et al., 2024), and VHELM supplies a multi-dimensional benchmark covering bias, fairness, toxicity and safety (Lee et al., 2024). Janghorbani and De Melo (2023) introduce a unified framework for stereotypical bias across modalities, while Raj et al. (2024) expose hidden biased associations that evade existing diagnostics.
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT-style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2022). Ghate et al. (2025) show that intrinsic bias is largely predictable from pre-training data, not architecture. Liu et al. (2024) offer a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases . Our controlled study extends this line by showing that size, scale, and loss composition interact non-linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five-fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.

### v4

# A Comprehensive Social Bias Audit of Contrastive Vision–Language Models

@@ Abstract
A Comprehensive Social Bias Audit of Contrastive Vision–Language Models
@@ 1 Introduction
Existing audits seldom answer this question because they vary only one factor or inspect a single model family, leaving the effects of model capacity, data volume, and data composition entangled. In language-only settings, scaling can either help or hurt fairness depending on objective and metric (Liang et al., 2021), but systematic evidence for VLMs is still limited.
@@ 2.2 Bias-Measurement Frameworks and Mitigation
Task-agnostic toolkits now quantify social bias in VLM outputs. MMBias measures stereotype leakage in captioning and VQA (Janghorbani and de Melo, 2023), whereas DebiasCLIP prepends learned visual tokens to attenuate gender stereotypes (Berg et al., 2022). Embedding-space analyses reveal that CLIP encodes the family $\leftrightarrow$ career gender stereotype (Bianchi et al., 2023; Liang et al., 2021). Prompt-level interventions, such as the Debiasing and VisDebiasing prefixes in ModSCAN, reduce but do not eradicate skew (Jiang et al., 2024). Beyond CLIP, BendVLM performs nonlinear, test-time debiasing of VLM embeddings without fine-tuning (Gerych et al., 2024), Association-Free Diffusion mitigates object-people stereotype transfer in text-to-image generation (Zhou et al., 2024), FairCoT leverages chain-of-thought reasoning in multimodal LLMs to iteratively refine prompts and improve fairness in text-to-image generation (Al Sahili et al., 2024), and VHELM supplies a multi-dimensional benchmark covering bias, fairness, toxicity and safety (Lee et al., 2024). Janghorbani and De Melo (2023) introduce a unified framework for stereotypical bias across modalities, while Raj et al. (2024) expose hidden biased associations that evade existing diagnostics.
@@ 2.3 Scaling Effects on Bias
Scaling model parameters or data volume is often viewed as a route to robustness, yet empirical findings remain mixed. In language models, larger GPT-style systems sometimes reduce toxic generation (Hoffmann et al., 2022) but can entrench occupational stereotypes (Liang et al., 2021). Ghate et al. (2025) show that intrinsic bias is largely predictable from pre-training data, not architecture. Liu et al. (2024) offer a probabilistic framework revealing that stereotype volatility increases with model size. For VLMs, Birhane et al. (2023) warned that expanding LAION from 400M to 5B images risks amplifying existing harms, whereas higher image resolution can mitigate certain gender biases . Our controlled study extends this line by showing that size, scale, and loss composition interact non-linearly: the same parameter increase that softens bias in CLIP amplifies it in OpenCLIP, and a five-fold data increase leaves CLIP’s skew almost unchanged while doubling that of OpenCLIP.